OpenStudy (goformit100):

Verify by the method of contradiction. p: root7 is irrational

OpenStudy (goformit100):

@Noemi95 @terenzreignz

terenzreignz (terenzreignz):

Yikes.... I think we need a new statement for this, a lemma... Let a be an integer... then if \(a^2\) is divisible by 7, then \(a\) is also divisible by 7.

OpenStudy (anonymous):

@terenzreignz 's got it gofor! (:

terenzreignz (terenzreignz):

We have to prove that, first... So, let a be an integer that is NOT divisible by 7 (proof by contrapositive) Then \[\large a = 7q + r \quad\quad\quad q,r\in\mathbb{Z}, \quad r<7\] then \[\large a^2=(7q+r)^2=49q^2+14qr+r^2=7(\color{blue}{7q^2+2qr})+r^2\]which is also not divisible by 7 since \(r^2 \) is not divisible by 7.

OpenStudy (goformit100):

Thanks

terenzreignz (terenzreignz):

thus, if \(a^2\) is divisible by 7, then a is also divisible by 7. That's a nice tool for this question... should we move to your next?